# TorchGDM: A GPU-Accelerated Python Toolkit for Multi-Scale Electromagnetic Scattering with Automatic Differentiation

### v2

@@ Abstract
We present “torchGDM”, a numerical framework for nano-optical simulations based on the Green’s Dyadic Method (GDM). This toolkit combines a hybrid approach, allowing for both fully discretized nano-structures and structures approximated by sets of effective electric and magnetic dipoles. It supports simulations in three dimensions and for infinitely long, two-dimensional structures. This capability is particularly suited for multi-scale modeling, enabling accurate near-field calculations within or around a discretized structure embedded in a complex environment of scatterers represented by effective models. Importantly, torchGDM is entirely implemented in PyTorch, a well-optimized and GPU-enabled automatic differentiation framework. This allows for the efficient calculation of exact derivatives of any simulated observable with respect to various inputs, including positions, wavelengths or permittivity, but also intermediate parameters like Green’s tensors. We anticipate that this toolkit will be valuable for applications merging nano-photonics and machine learning, as well as for solving nano-photonic optimization and inverse problems, such as the global design and characterization of metasurfaces, where optical interactions between structures are critical.
@@ I Introduction
Gustav Mie expanded the fields scattered by spherical particles in multipolar contributions using spherical harmonics.mieBeitrageZurOptik1908 Such multipole decomposition of electromagnetic fields is today also frequently used for particles of non-spherical geometries,alaeeElectromagneticMultipoleExpansion2018 ; alaeeExactMultipolarDecompositions2019 ; majorelGeneralizingExactMultipole2022 ; kildishevArtFindingOptimal2025 which is the theoretical foundation of the T-matrix method. watermanMatrixFormulationElectromagnetic1965 ; petersonMatrixElectromagneticScattering1973 ; mishchenkoTmatrixMethodIts2010 ; egelSMUTHIPythonPackage2021 ; theobaldSimulationLightScattering2021 ; schebarchovMultipleScatteringLight2022 Likewise, effective electric (ED) and magnetic dipole (MD) point-polarizabilities are a commonly used semi-analytical model, appealing thanks to its formal simplicity. It can be used to describe large assemblies of nano-scatterers, in cases where the truncation of the multipole expansion at the dipole order is valid.mulhollandLightScatteringAgglomerates1994 ; chaumetCoupleddipoleMethodMagnetic2009 ; patouxPolarizabilitiesComplexIndividual2020 However, these methods allow a description only outside of a circumscribing sphere around non-spherical particles. Attempts to overcome this problem only offer partial solutions to the problem.martinTmatrixMethodClosely2019
@@
Therefore, semi-analytical methods like the T-matrix approach imply limitations for certain use-cases. Examples are dense arrays of non-spherical particles where the circumscribing spheres overlap, or scenarios with illumination by local emitters like fluorescent molecules or quantum dots close to the particles or inside the circumscribing sphere, or when internal fields are required. For perfect periodic systems, local emitters also pose a challenge. A phased-array scanning method can be used to decompose the dipole field in periodic contributions, but also then, the local light source needs to be sufficiently far from the array.capolinoComparisonMethodsCalculating2007 ; lunnemannLocalDensityOptical2016
@@
Recently proposed methods like the global polarizability matrix method (“GPM”)bertrandGlobalPolarizabilityMatrix2020 or the topological skeleton of multiple-multipoleslamprianidisTranscendingRayleighHypothesis2023 distribute several multipolar sources across the particle. However, such models remain approximative at least within the volume of the nanostructures.
@@
To overcome these limitations, we propose to combine the semi-analytical “GPM” method, based on electric and magnetic dipole pairs, with a volume integral approach, the Green’s Dyadic Method (GDM).martinGeneralizedFieldPropagator1995 ; girardFieldsNanostructures2005 In our description, particles can be either represented by sets of effective electric and magnetic point dipole pairs (GPM),munDescribingMetaAtomsUsing2020 or as a volume discretization (in 2D: surface discretization). The latter is strongly inspired by our former GDM implementation “pyGDM”wiechaPyGDMPythonToolkit2018 ; wiechaPyGDMNewFunctionalities2022 . Dipole emitters can be used as illumination source, and the photonic local density of states (LDOS) and Green’s tensors in complex environments can be calculated.cazeSpatialCoherenceComplex2013 ; carminatiElectromagneticDensityStates2015
@@
The first peculiarity of our “TorchGDM” toolkit is that both types of particle representations (effective polarizability models and volume discretizations) can be mixed in a single multi-scale simulation. The second particularity is that torchGDM is fully implemented in the automatic differentiation (autodiff or “AD”) framework PyTorch, making any simulation entirely differentiable with respect to any input (like positions, wavelength, permittivity, …) or intermediate variables (e.g. Green’s tensors).
@@ I.1 Similar software
Before we describe torchGDM in detail, we want to provide a non-exhaustive selection of openly available electrodynamics simulation toolkits and briefly summarize their specific strengths.
@@ I.1.1 Surface integral method
Best for single particle simulations.
@@ I.1.2 Discrete Dipole Approximation - iterative solver
Good for single particle simulations.
@@ I.1.3 Discrete Dipole Approximation - direct solver
Good for single particle, multi-illumination simulations. Smaller particle sizes and reduced accuracy, due to discretization limitations.
@@ I.1.4 T-matrix method
Best for large multi particle simulations. Single particle t-matrices need to be calculated beforehand.
@@ I.1.5 RCWA
Best for periodic structures, possibly in layered systems.
@@ I.1.6 FDTD
Very general method. Well suited for single particles and periodic simulations, as well as for larger models thanks to a friendly scaling behavior.
@@ I.1.7 This work
The here presented tool “TorchGDM” implements a hybrid coupled dipole method that allows to combine volume discretizations as in the GDM,girardFieldsNanostructures2005 with effective point source models, like effective electric/magnetic dipole polarizabilitiesmunDescribingMetaAtomsUsing2020 or global polarizability matrices (GPMs).bertrandGlobalPolarizabilityMatrix2020 The latter allow efficient and accurate modelling of near-fields also inside the circumscribing sphere of non-spherical particles. TorchGDM supports 3D as well as 2D simulations. The code is entirely written using “PyTorch” as backend, an automatic differentiation framework, which is particularly popular in the deep learning community. This allows to calculate derivatives of any possible observable with respect to arbitrary input or intermediate parameters. Finally, PyTorch offers full GPU support, simulations can therefore be run on accelerator hardware with no additional effort.
Figure 1: (a) Reference system and angular conventions in torchGDM. (b) Sketch of the global polarizability matrix (GPM) effective model formalism.bertrandGlobalPolarizabilityMatrix2020 A GPM describes a particle through a set of local (dark blue) and non-local polarizabilities.This means, the incident field at one specific polarizability location induces electric and magnetic dipole moments at all polarizability locations in the GPM. (c-d) available structure types for (c) 3D and (d) 2D simulations. In addition to discretized structures, effective dipole models are supported that may contain several non-local dipoles (global polarizability matrix, “GPM” (b)), or a single pair of an electric and a magnetic dipole (“dp pair”). Effective models can be extracted from discretized structures, T-Matrices or from Mie theory.
II Formalism
@@
$^{\prime}.\end{multlined}\mathbf{F}(\mathbf{r})=\mathbf{F}_{0}(\mathbf{r})+\int_{V}\boldsymbol{\mathbbm{G}}(\mathbf{r},\,\mathbf{r}^{\prime})\bm{\chi}(\mathbf{r}^{\prime})\mathbf{F}(\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime}.$ (1)
@@
$\,.\end{multlined}\mathbf{F}(\mathbf{r}_{i})=\mathbf{F}_{0}(\mathbf{r}_{i})+\sum_{j=1}^{N}\boldsymbol{\mathbbm{G}}(\mathbf{r}_{i},\,\mathbf{r}_{j})\bm{\alpha}_{j}\mathbf{F}(\mathbf{r}_{j})\,.$ (5)
@@
$\end{multlined}\mathbf{F}_{0}(\mathbf{r}_{i})=\sum_{j=1}^{N}\mathbf{F}(\mathbf{r}_{j})\Big{[}\delta_{ij}\mathbf{I}-\boldsymbol{\mathbbm{G}}(\mathbf{r}_{i},\,\mathbf{r}_{j})\bm{\alpha}_{j}\mathbf{F}(\mathbf{r}_{j})\Big{]}\,,$ (7)
@@
Note that it has been demonstrated in recent literature, that the solution of coupled systems in linear optics can be efficiently performed with automatic differentiation.delhougneLearnedIntegratedSensing2020 ; gargInverseDesignedDispersiveTimeVarying2025
@@ II.2 Extraction of effective models
Please note that a warning is triggered if the $L_{2}$ residual is larger than an empirically found threshold (which can also be manually set by the user). In this case, the structure is probably too large and different illumination directions lead to different dipole moments, which cannot be represented in the model. To solve such issues, one should increase the number of effective dipoles in the model. Be aware that this only tests the residuals of the minimization problem. The model may still work badly with other illuminations than used during the extraction. To this end, the accuracy of effective models can be automatically tested in a small scattering simulation, by passing the parameter test_accuracy=True to the extraction function. This will provide a relative error with respect to cross sections and near-fields obtained from a full simulation (or a T-Matrix calculation), estimating the fidelity of the effective model approximation.
@@ II.3 Effective polarizabilities of spherical and cylindrical core-shell particles and T-Matrices
TorchGDM contains also a tool to obtain effective polarizability models for spherical and cylindrical core-shell particles as well as for arbitrary 3D and 2D T-Matrices in a homogeneous environment of real refractive index $n_{\text{env}}$. GPM models of Mie spheres, cylinders and T-Matrices are obtained using the scattered field matching technique described above. The model for a single pair of point dipoles can be obtained using the Mie scattering coefficients $a_{n}$ and $b_{n}$ (see figure 1c-d).bohrenAbsorptionScatteringLight1998 ; garcia-etxarriStrongMagneticResponse2011 In the 3D case, we get:
$\begin{split}\bm{\alpha_{pE}}&=\mathbf{I}\frac{3i}{2}\ \frac{a_{1}}{k_{0}^{3}n_{\text{env}}}\\ \bm{\alpha_{mH}}&=\mathbf{I}\frac{3i}{2}\ \frac{b_{1}}{k_{0}^{3}n_{\text{env}}^{3}}\,,\end{split}$ (13)
@@
$\begin{split}\alpha_{2D,pE}^{TE}&=\frac{i}{\pi}\ \frac{a_{1}}{k_{0}^{2}}\\ \alpha_{2D,mH}^{TE}&=\frac{2i}{\pi}\ \frac{b_{1}}{k_{0}^{2}n_{\text{env}}^{2}}\,,\end{split}$ (14)
@@ II.4 Hybrid discretization GDM
Figure 3: (a) Example 3D geometry. A discretized structure $V$ with $N$ cubic volume elements $v_{i}$ (orange) is coupled to two other structures (green and blue), represented each by a pair of electric-electric and magnetic-magnetic effective polarizabilities $\bm{\alpha}_{\text{eff}1}$ and $\bm{\alpha}_{\text{eff}2}$. (b) 2D projection of the geometry. (c) Mixed polarizability linear coupled system, as solved in torchGDM (by inversion of the coupling matrix).
As mentioned before, torchGDM allows to arbitrarily combine structures approximated by effective point polarizability models and volume discretized particles. In practice, we write our system of coupled equations (5) by separating all electric and all magnetic fields, as illustrated in figure 3. Technically this facilitates removal of rows and columns of zero magnetic polarizability from the system of equations, to reduce the memory footprint and computational cost for inversion.
@@ II.5 Observables
  • Field gradients.
  • Total extinction, absorption and scattering cross sections.
  • Total and differential scattered intensity via repropagation (more accurate for very small absorptive particles).
@@ II.6 Implementation technical details
  • The total scattering and absorption cross sections of GPM effective models cannot be calculated with the optical theorem right now, only the extinction cross section works correctly. This is a consequence of the non-local character of the effective dipoles in this case (each dipole moment is a function of the illumination fields at at all other dipoles’ locations). To obtain the scattering cross section, integration of the farfield scattering on a spherical surface is necessary. Absorption can then be obtained as $\sigma_{\text{abs}}=\sigma_{\text{ext}}-\sigma_{\text{sca,ff}}$.
  • Tabulated materials permittivities are supported using the “yaml” format from https://refractiveindex.info/.polyanskiyRefractiveindexinfoDatabaseOptical2024 Either tabulated data or Sellmeier models are supported so far. The permittivity data is bi-linearly interpolated between tabulated wavelengths.
  • 2D Green’s tensors: Due to current PyTorch limitations, Hankel functions are evaluated using a combination of Bessel functions and recurrence relations. A technical consequence is that only integer order and pure real or pure imaginary arguments are supported. This can be limiting, for example in applications that require integration along complex frequencies.
@@
  • Field gradients require differentiation of many output values with respect to many input values (field components at multiple positions). Autograd is not efficient in such cases (neither backward mode nor forward mode), therefore TorchGDM’s field gradient routine is based on numerical center differences. Field gradient calculation may of course be also implemented using PyTorch autodiff. This will be generally slow due to the requirement of building many computational graphs, but it may be interesting in some cases, for example to evaluate higher order derivatives with high accuracy.
@@
III Implementation structure and API
We present in the following the basic package structure and the most relevant elements of the interface. For a full technical documentation we refer the reader to the online documentation.
III.1 Package structure

### v3

@@ Abstract
We present “torchGDM”, a numerical framework for nano-optical simulations based on the Green’s Dyadic Method (GDM). This toolkit combines a hybrid approach, allowing for both fully discretized nano-structures and structures approximated by sets of effective electric and magnetic dipoles. It supports simulations in three dimensions and for infinitely long, two-dimensional structures. This capability is particularly suited for multi-scale modeling, enabling accurate near-field calculations within or around a discretized structure embedded in a complex environment of scatterers represented by effective models. Importantly, torchGDM is entirely implemented in PyTorch, a well-optimized and GPU-enabled automatic differentiation framework. This allows for the efficient calculation of exact derivatives of any simulated observable with respect to various inputs, including positions, wavelengths or permittivity, but also intermediate parameters like Green’s tensor components, which can be interesting for physics informed deep learning applications. We anticipate that this toolkit will be valuable for applications merging nano-photonics and machine learning, as well as for solving nano-photonic optimization and inverse problems, such as the global design and characterization of metasurfaces, where optical interactions between structures are critical.
@@ I Introduction
Gustav Mie expanded the fields scattered by spherical particles in multipolar contributions using spherical harmonics.mieBeitrageZurOptik1908 Such multipole decomposition of electromagnetic fields is today also frequently used for particles of non-spherical geometries,alaeeElectromagneticMultipoleExpansion2018 ; alaeeExactMultipolarDecompositions2019 ; majorelGeneralizingExactMultipole2022 ; kildishevArtFindingOptimal2025 which is the theoretical foundation of the T-matrix method. watermanMatrixFormulationElectromagnetic1965 ; petersonMatrixElectromagneticScattering1973 ; mishchenkoTmatrixMethodIts2010 ; egelSMUTHIPythonPackage2021 ; theobaldSimulationLightScattering2021 ; schebarchovMultipleScatteringLight2022 Likewise, effective electric (ED) and magnetic dipole (MD) point-polarizabilities are a commonly used semi-analytical model, appealing thanks to its formal simplicity. It can be used to describe large assemblies of nano-scatterers, in cases where the truncation of the multipole expansion at the dipole order is valid.mulhollandLightScatteringAgglomerates1994 ; chaumetCoupleddipoleMethodMagnetic2009 ; patouxPolarizabilitiesComplexIndividual2020 ; fradkinNanoparticleLatticesBases2020 However, these methods allow a description only outside of a circumscribing sphere around non-spherical particles. Earlier attempts to overcome this problem only offered partial solutions to the problem.martinTmatrixMethodClosely2019
@@
Therefore, semi-analytical methods like the T-matrix approach imply limitations for certain use-cases. Examples are dense arrays of non-spherical particles where the circumscribing spheres overlap, or scenarios with illumination by local emitters like fluorescent molecules or quantum dots close to the particles or inside the circumscribing sphere, or when internal fields are required. For perfect periodic systems, local emitters also pose a challenge. A phased-array scanning method can be used to decompose the dipole field in periodic contributions, however this technique is algorithmically and computationally very demanding.capolinoComparisonMethodsCalculating2007 ; lunnemannLocalDensityOptical2016 ; kogonFDTDModelingPeriodic2020
@@
Recently proposed methods like the global polarizability matrix method (“GPM”)bertrandGlobalPolarizabilityMatrix2020 or the topological skeleton of multiple-multipoleslamprianidisTranscendingRayleighHypothesis2023 distribute several multipolar sources across the particle. However, at least inside the volume of the nanostructures such models remain approximative.
@@
To overcome these limitations, we propose to combine the semi-analytical “GPM” method with a volume integral approach, the Green’s Dyadic Method (GDM).martinGeneralizedFieldPropagator1995 ; girardFieldsNanostructures2005 In our description, particles can be either represented by sets of effective electric and magnetic point dipole pairs (GPM),munDescribingMetaAtomsUsing2020 or as a volume discretization (in 2D: surface discretization). The latter is strongly inspired by our former GDM implementation “pyGDM”wiechaPyGDMPythonToolkit2018 ; wiechaPyGDMNewFunctionalities2022 . Dipole emitters can be used as illumination source, and the photonic local density of states (LDOS) and Green’s tensors in complex environments can be calculated.cazeSpatialCoherenceComplex2013 ; carminatiElectromagneticDensityStates2015
@@
The first peculiarity of our “TorchGDM” toolkit is that both types of particle representations (effective polarizability models and volume discretizations) can be mixed in a single multi-scale simulation. The second particularity is that torchGDM is fully implemented in the automatic differentiation (autodiff or “AD”) framework PyTorch, making any simulation entirely differentiable with respect to any input (like positions, wavelength, permittivity, …) or intermediate variables (e.g. Green’s tensor components).
@@ I.1 Similar software
Before we describe torchGDM in detail, we want to give a non-exhaustive selection of electrodynamics simulation methods, briefly discuss their pros and cons, and provide a selection of openly available toolkits.
@@ I.1.1 Surface integral method
Best for single particle simulations, can handle complex shapes. Possibility to implement solver schemes with guaranteed convergence. Usually too computationally heavy for many-particle simulations.
@@ I.1.2 Discrete Dipole Approximation - iterative solver
Good for single particle simulations. Convergence is not guaranteed and it can show unexpected convergence problems. It is possible to improve the accuracy for instance using filtering techniques pillerIncreasingPerformanceCoupleddipole1998 . Improvements have also been suggested for specific applications such as high aspect ratio particles smunevRectangularDipolesDiscrete2015 , high index dielectric materials chaumetCoupledDipoleMethod2004 , or periodic structures chaumetGeneralizationCoupledDipole2003 . With the iterative solver, each incident field configuration needs an individual simulation run, therefore it can be expensive for multiple illuminations.
@@ I.1.3 Discrete Dipole Approximation - direct solver
Good for single particle. In principle, same as iterative DDA, but limited to smaller particle sizes and reduced accuracy, due to discretization limitations. Very well suited for multi-illumination simulations, because the inverted coupling system can be reused. Simple to implement.
@@ I.1.4 T-matrix method
Best for large multi particle simulations. Only solves the coupling between particles. Single particle t-matrices need to be calculated beforehand. Particles must not be closer than their circumscribing spheres. To the best of our knowledge, so far no automatic differentiation implementation exists.
@@ I.1.5 RCWA
Fourier modal method. Best for periodic structures, possibly in layered systems. Not ideal for single particles due to implicit periodic assumption. Very efficient in 1D and 2D, as well as for simple 3D structures. Expensive for complex 3D structures. Can show convergence problems in structures with complex interfaces (like dielectric/metal).
@@ I.1.6 FDTD
Very general method. Well suited for single particles and periodic simulations, as well as for larger models thanks to a friendly scaling behavior. Evanescent fields or strong gradients require fine meshes and can pose convergence problems. A simple FDTD code is easy to implement.
@@ I.1.7 This work
The here presented tool “TorchGDM” implements a hybrid coupled dipole method that allows to combine volume discretizations as in the discrete dipole approximation (in the following we will call it the Green’s Dyadic Method, GDMgirardFieldsNanostructures2005 ), with effective point source models, like effective electric/magnetic dipole polarizabilitiesmunDescribingMetaAtomsUsing2020 and global polarizability matrices (GPMs).bertrandGlobalPolarizabilityMatrix2020 GPMs allow efficient and accurate modelling of large particles and calculation of near-fields also inside the circumscribing sphere of non-spherical particles. TorchGDM supports 3D as well as 2D simulations. With the exception of an interface to the “treams” T-Matrix toolkit, the code is entirely written using “PyTorch” as backend, a popular automatic differentiation framework from the deep learning community. This allows to calculate derivatives of any possible observable with respect to arbitrary input or intermediate parameters. Finally, PyTorch offers full GPU support, simulations can therefore be run on accelerator hardware with no additional effort.
The most important capabilities are:
  • 2D and 3D simulations
  • Volume discretized particle models
  • Effective dipoles based particle models
  • Overlapping circumscribing spheres of effective models
  • Combining particle types for multi-scale simulations
  • Various far-field observables (complex far-field, total and differential cross sections, radiation patterns, exact multipole decomposition)
  • Various near-field observables (electric and magnetic near-fields, fields inside discretized structures, chirality, Poynting vector, energy flux, field gradients, LDOS, Green’s tensors, multipole decomposition)
  • Automatic differentiation of any simulation calculation (with very few exceptions, see section II.7)
  • Various visualization tools
  • Interfaces to external T-Matrix and Mie tool (“treams”) for effective model creation (not AD compatible)
We believe that torchGDM will be particularly useful at medium size or multi-scale scattering problems, in cases where derivatives are required. Examples include design tasks like metasurface design, for instance when more than next-neighbor interactions are relevant for an accurate description. AD may also be useful in problems where a numerical model is to be fitted to experimental results. The differentiability can also be used to train neural networks within the simulation description, for example to learn effective Green’s tensors or structure models.
In the future, further automatic differentiable components may be added. Possible developments include differentiable Mie theory or T-Matrix support, or wavefront propagation via the angular spectrum method. Layered media or periodic structures would also be interesting future developments, that could be added through Green’s tensors as additional environment classes without the need to modify the main code itself. Finally, the full inversion scheme currently limits torchGDM to roughly 10000 coupled dipoles. This could be improved through iterative solver schemes. This could also be modularly implemented as alternative linear system class.
Figure 1: (a) Reference system and angular conventions in torchGDM. (b) Sketch of the global polarizability matrix (GPM) effective model formalism.bertrandGlobalPolarizabilityMatrix2020 A GPM describes a particle through a set of local (dark blue) and non-local polarizabilities.This means, the incident field at one specific polarizability location induces electric and magnetic dipole moments at all polarizability locations in the GPM. (c-d) available structure types for (c) 3D and (d) 2D simulations. In addition to discretized structures, effective dipole models are supported that may contain several non-local dipoles (global polarizability matrix, “GPM” (b)), or a single pair of an electric and a magnetic dipole (“dp pair”). Effective models can be extracted from discretized structures, T-Matrices or from Mie theory. Note that the Mie and T-Matrix extractions are not AD compatible.
II Formalism and implementation
@@
[…]
@@
[…]
@@
[…]
@@
Note that automatic differentiation of the LU decomposition algorithm is numerically stable, which has also been demonstrated in the context of the coupled dipoles formalism in recent literature.delhougneLearnedIntegratedSensing2020 ; gargInverseDesignedDispersiveTimeVarying2025
@@ II.2 Extraction of effective models
Please note that a warning is triggered if the $L_{2}$ residual is larger than an empirically found threshold (which can also be manually set by the user). In this case, the structure is probably too large and different illumination directions lead to different dipole moments, which cannot be represented in the model. To solve such issues, one should increase the number of effective dipoles for the particle model. Be aware that this only tests the residuals of the minimization problem. The model may still work badly with other illuminations than used during the extraction. Therefore, the accuracy of effective models should always be tested in a small scattering simulation. This can be done automatically, by passing the parameter test_accuracy=True to the extraction function. This will provide a relative error with respect to cross sections and near-fields obtained from a full simulation (or a T-Matrix calculation), estimating the fidelity of the effective model approximation.
@@ II.3 Effective polarizabilities of spherical and cylindrical core-shell particles and T-Matrices
TorchGDM contains also a tool to obtain effective polarizability models for spherical and cylindrical core-shell particles as well as for arbitrary 3D and 2D T-Matrices in a homogeneous environment of real refractive index $n_{\text{env}}$. GPM models of Mie spheres, cylinders and T-Matrices are obtained using the scattered field matching technique described above. It is possible to use other Maxwell solvers to extract an effective model, either through the simulated illumination and scattered fields, or via a detour by first extracting the T-Matrix of a nanostructure asadovaTmatrixRepresentationOptical2025 , and convert this through TorchGDM’s interface to “treams” into a global polarizability matrix effective model.
The model for a single pair of point dipoles can be obtained using the Mie scattering coefficients $a_{n}$ and $b_{n}$ (see figure 1c-d).bohrenAbsorptionScatteringLight1998 ; garcia-etxarriStrongMagneticResponse2011 In the 3D case, we get:
[…]
@@
[…]
@@ II.4 Hybrid discretization GDM
Figure 3: (a) Example 3D geometry. A discretized structure $V$ with NN cubic volume elements $v_{i}$ (orange) is coupled to two other structures (green and blue), represented each by a pair of electric-electric and magnetic-magnetic effective polarizabilities $\bm{\alpha}_{\text{eff}1}$ and $\bm{\alpha}_{\text{eff}2}$. (b) 2D projection of the geometry. (c) Mixed polarizability linear coupled system, as solved in torchGDM (by inversion of the coupling matrix).
As mentioned before, torchGDM allows to arbitrarily combine structures approximated by effective point polarizability models and volume discretizations. In practice, we write our system of coupled equations (5) by separating all electric and all magnetic fields, as illustrated in figure 3. Technically this facilitates removal of rows and columns of zero magnetic polarizability from the system of equations, to reduce the memory footprint and computational cost for inversion.
@@ II.5 Observables
  • Field gradients (via automatic differentiation).
  • Total extinction, absorption and scattering cross sections.markelExtinctionScatteringAbsorption2019
  • Total and differential scattered intensity via repropagation (more accurate for very small absorptive particles).chaumetCoupleddipoleMethodMagnetic2009
@@ II.6 Implementation technical details
  • The total scattering and absorption cross sections of GPM effective models can right now not be calculated with the optical theorem, only the extinction cross section works correctly with GPM models (models with more that 1 effective dipole pair). The effective dipoles in these GPMs are non-local: each dipole moment is a function of the illumination fields also at all other dipoles’ locations. This non-local response renders the computation of the work that the field exerts on the effective dipoles complicated, and we have not implemented this so far. Currently, to obtain the scattering cross section, integration of the farfield scattering on a spherical surface is necessary. Absorption can then be obtained as $\sigma_{\text{abs}}=\sigma_{\text{ext}}-\sigma_{\text{sca,ff}}$.
  • Materials permittivities from literature are supported using the “yaml” format from https://refractiveindex.info/.polyanskiyRefractiveindexinfoDatabaseOptical2024 Either tabulated data or Sellmeier models are supported so far. The permittivity data is bi-linearly interpolated between tabulated wavelengths.
  • 2D Green’s tensors: Due to current PyTorch limitations, Hankel functions are evaluated using a combination of Bessel functions and recurrence relations. A technical consequence is that only integer order and purely real arguments are supported so far. This can be limiting for example in applications that require integration along complex frequencies in 2D simulations.
@@
  • Field gradients require differentiation of many output values with respect to many input values (field components at multiple positions). Autograd is typically not efficient in such cases (neither backward mode nor forward mode). However, PyTorch offers a composable functions transforms module “torch.func”, which allows very efficient calculation of jacobians through automatic vectorization. Since torch.func is still in beta, torchgdm implements as a fallback alternative field gradients based on finite differences. Both methods yield similar accuracy and are automatically differentiable. Field gradient calculation may of course be also implemented using PyTorch autodiff. This will be generally slow due to the requirement of building many computational graphs, but it may be interesting in some cases, for example to evaluate higher order derivatives with high accuracy.
@@
II.7 Automatic differentiation limitations
TorchGDM is entirely written in PyTorch, but some tools use external libraries or act as interfaces to third party software. In consequence, these functions are not compatible with automatic differentiation. We give here a list of the functionalities that are not fully torch AD-capable:
  • Mie theory: All tools that use Mie theory interface with “treams”beutelTreamsTmatrixbasedScattering2024 and thus do not support AD. For a simulation workflow this means that Mie scattering sections are not AD-capable at all. These tools are meant to serve as a comparison or benchmark baseline. Mie-based core-shell effective models can be included in an autodiff workflow after their non-AD capable initialization.
  • T-matrix conversion also uses “treams”beutelTreamsTmatrixbasedScattering2024 and does not support AD. T-Matrix based effective structure models can be included in an autodiff workflow after their non-AD capable initialization.
  • In the case when only an integer number of effective dipoles is given for GPM conversion, torchGDM uses spectral clustering from the “scikit-learn”pedregosaScikitlearnMachineLearning2011 library to distribute effective sources inside the particle geometry. This is not AD compatible. The user can specify the effective dipole locations manually as a list of positions. In this case, GPM extraction is autodifferentiable.
  • It is possible to load an image and convert it into a torchGDM discretized structure. However, the function “tg.struct3d.from_image” uses “PIL” to load and pre-process the image file, and therefore this operation is not autodiff capable.